# AIReg-Bench: Benchmarking Language Models That Assess AI Regulation Compliance

### v1

@@ 5.2 LLMs for legal, compliance, and AIR compliance tasks
Researchers have applied LLMs to a wide variety of legal tasks (Ma et al., 2024; Lai et al., 2024; Siino et al., 2025). This includes various compliance assessment tasks (Hassani, 2024; Bolton et al., 2025; Chen et al., 2024; Wang et al., 2025a), including AIR compliance assessments. For example, Makovec et al. (2024) input datasets, model cards, README files, or other AI project artifacts into a RAG-enhanced GPT-4 that accesses relevant portions of the AIA to predict the compliance level of the AI system depicted in the input. Davvetas et al. (2025) use a RAG-equipped LLM (mistral-small3.2) that takes, as input, certain features of an AI system (such as the type of AI system and the intended use) and outputs the risk-level of the AI system according to the AIA. Similarly, Kővári et al. (2025) use in-context learning and RAG to create a chatbot that can help users self-assess compliance with the AIA. Meanwhile, Li et al. (2025) test various LLMs’ ability to act as a rudimentary AIA compliance checker by accepting a hypothetical AI system as context and predicting whether it is prohibited by, permitted by, or out of scope of the AIA. A series of interlinked studies by Nokia Bell Labs employ LLMs to support AI practitioners in AIR compliance subtasks such as populating impact assessment reports (Bogucka et al., 2024a; Herdel et al., 2024; Bogucka et al., 2024b). Sovrano et al. (2025) use an LLM to help human-drafted technical documentation align with the requirements of AIA Article 11.
@@ 5.3 The EU AI Act
The AIA went into force in August 2024 (Lomas, 2024) and sets forth harmonized requirements for AI systems and models placed on the market or put into service in the EU (EU, 2024, Art. 1-2). In laying out requirements for these AI systems, the AIA leverages a “risk-based” approach (Mahler, 2022), by which the exact requirements that apply to a system are a function of its perceived degree of risk. Here, the most demanding requirements are reserved for those AI systems deemed to be high-risk (EU, 2024, Art. 6). Such high-risk AI systems (HRAI) must satisfy a number of requirements (EU, 2024, Chap. III, Sec. 2). Among those, the requirements that we have made the focus of our benchmark dataset relate to risk management systems, data and data governance, record keeping, human oversight, as well as accuracy, robustness, and cybersecurity (EU, 2024, Art. 9, 10, 12, 14, 15).
@@ 6.1 Challenges of legal benchmarking
Compliance assessments are subjective. Complicating legal benchmarking is the fact that legal tasks often involve subjective judgments (Guha et al., 2024; Ma et al., 2023). Compliance, in particular, has been described as “not binary” (Wu & van Rooij, 2021). This was one motivation for our use of Likert scale annotations keyed to the “probability” of compliance rather than binary labels of “compliant” or “non-compliant.” The Likert scale is viewed as a reliable way to “transform subjective qualitative data into quantifiable metrics” (Koo & Yang, 2025) and has previously been used for benchmarking of LLMs in subjective realms (Bojić et al., 2025).
@@
Compliance assessments are a moving target. It has been said that compliance assessments require clear and specific guidelines, including relevant case law (Kilian et al., 2025; Schuett, 2024). But, in its present state, the AIA arguably lacks these. The text of the law has not been interpreted by courts (Yew et al., 2025). The obligations outlined in the AIA have yet to be clarified by accompanying technical standards (European Committee for Electrotechnical Standardization, 2024; European Commission, 2022). They are also subject to ongoing amendments (EU, 2024, Art. 96). Accordingly, AIReg-Bench should be viewed as a snapshot of AIA conformity assessments in September 2025. It does not and cannot reflect developments occurring after this date.
@@ 6.2 Challenges of LLM-driven sample generation
Benchmark dataset samples should be representative of real-world data (Sourlos et al., 2024). Broadly speaking, there is evidence that LLMs (especially larger ones) can effectively produce synthetic samples satisfying this criteria (Maheshwari et al., 2024). More directly relevant here, it has been shown that LLMs can effectively generate (or improve) legal documents (Su et al., 2025; Lin & Cheng, 2024; Hemrajani, 2025; Gray et al., 2025), technical specifications (Xie et al., 2025), compliance documentation (Wang et al., 2025d; Hassani, 2024; Kumar & Roussinov, 2024), and even the type of technical documentation required under the AIA (Sovrano et al., 2025).
However, some are sceptical about the ability of LLMs to generate realistic outputs in these domains (Posner & Saran, 2025; Roberts, ; Shen et al., 2022). Critics point out that LLMs often lack domain-specific tacit knowledge, have difficulty maintaining coherent reasoning across extended contexts, and may hallucinate facts or references (Rasiah et al., 2024; Huang et al., 2024; Dahl et al., 2024; Magesh et al., 2025). Regarding legal tasks specifically, critics note how LLMs’ struggle to interpret legal terminology, grasp case context, and execute complex analyses, potentially resulting in errors (Wang et al., 2024; Roberts, ; Shen et al., 2022). With this in mind, a number of guardrails were put in place to enforce plausibility in our dataset samples and to create transparency around whether and where those guardrails fell short. For example:
  • The sample generation method was informed by the series of interviews with actual compliance assessment experts, described in Appendix D. We also consulted similar interviews performed by Li & Goel (2025) and the recommended protocols for manual compliance assessments for the AIA (Floridi et al., 2022; Thelisson & Verma, 2024; Lillo Campoy et al., 2024; Palumbo et al., 2025) and other AIR (The Institute of Internal Auditors, 2023; National Institute of Standards and Technology, 2023; Brogle et al., 2025).
@@ 6.3 Challenges of LLM-driven compliance analyses
Beyond the challenges of generating realistic documentation samples with LLMs, there may also be hurdles to using LLMs to perform legal analyses on text. Although some research efforts have found that LLMs match or exceed human accuracy when performing such analyses (Martin et al., 2024), others have questioned whether LLMs can perform this task effectively (Buckland, 2023; Doyle & Tucker, 2024; Li et al., 2024; Network, 2025; Mik, 2024). While the evidence presented in this paper is not definitive, we hope that our benchmark offers an initial step towards clarifying how well LLMs perform at compliance assessments, both in comparison to human experts and to one another. Our hope is that this will inspire other efforts to quantitatively evaluate the performance of LLMs at AIR compliance analyses and other legal tasks.
@@ 6.4 Risks of LLM-driven compliance analyses
As discussed in Section 1, there are potential benefits to LLM-driven AIR compliance analyses. However, it is important to point out that they could also carry risks. For example, given LLMs’ ongoing tendency to make errors, some argue that there may be dire consequences when lawyers place “blind faith in an LLM” (Moriarty, 2023) and that it may even represent a violation of professional ethics (Browning, 2024). To help inform this conversation, and avoid generalization, we believe it is invaluable to quantitatively evaluate the performance of LLMs for the task at hand, comparing their performance to human performance at that task in an evidence-based manner. Hence this benchmark.
@@ 7 Future Work
This benchmark should serve as a starting point for tracking LLM progress at AIR compliance assessments, rather than as a finish line indicating readiness for deployment in legal practice. Moving towards that goal will require additional benchmarks that build upon and extend AIReg-Bench. Some suggested directions for extensions to AIReg-Bench are outlined below.
@@
Extension to further LLM-powered annotators. It would be valuable to extend our benchmarking in Section 4 to include more models: including fine-tuned legal LLMs, which some argue perform better at legal tasks (Fei et al., 2023; Dominguez‑Olmedo et al., 2024), as well as LLMs with tool-use (e.g., incorporating RAG or web search) (Makovec et al., 2024; Davvetas et al., 2025; Wang & Yuan, 2025). Though, while to our knowledge, no such models exist yet, there would also be value in fine-tuning LLMs for AIA or AIR compliance and then evaluating them using AIReg-Bench.
Extension to real-world documentation. AIReg-Bench consists of LLM-generated excerpts of technical documentation for the AIA. We relied on these LLM-generated excerpts, whose plausibility we manually verified, since public access to real technical documentation is limited, perhaps due to the confidentiality or the relative nascency of the AIA. That said, we recognize the value, specifically as it relates to construct validity, of a benchmark built from real instances of AIA technical documentation. We therefore encourage any actors with access to such documentation to consider publishing (an anonymized version of) it.
Extension to AIA technical standards. Like other EU product regulation, the AIA will utilize harmonized standards, i.e., “more concrete” (Siegmann & Anderljung, 2022) technical specifications prepared by the EU’s external standardization organizations (CEN, CENELEC and ETSI). Compliance with these specifications, which are still under development (European Committee for Electrotechnical Standardization, 2024; European Commission, 2022), will “have the legal effect of establishing a presumption of conformity” with the AIA (Mazzini & Scalzo, 2023). Once issued, it would be important to pass these specifications, as additional context, into any AIA compliance assessment algorithm.
Extension to multi-turn interactions. Our interviews of AIR compliance professionals suggested that, in practice, compliance assessments often involve a long and complex dialogue between legal teams, technical staff, and regulators (see Appendix D). AIReg-Bench condenses the entire compliance assessment process into a single-turn interaction, based on a fixed set of synthetic artifacts. Such scoping and simplifications are common in benchmarking, though many scholar stress the need to shift towards more interactive modes of evaluating AI system capabilities (Ibrahim et al., 2024; Eriksson et al., 2025). Future benchmarks could build on AIReg-Bench by evaluating LLMs’ multi-turn ability to collaborate with human teams and contribute to complex legal dialogues (Kővári et al., 2025), rather than merely producing one-shot assessments.
@@ 8 Conclusion
Acknowledgments
We wish to acknowledge the Google Academic Research Awards for their support. Additionally, we wish to acknowledge the support of the OpenAI Researcher Access Program for providing subsidized API credits used in this research. We thank Bhagyashree Pancholy and Lara Groves for their insights during the early stages of this project.
Appendix A LLM usage in this work
LLMs were used in this work as follows:
  • LLMs were used to aid or polish writing. Specifically, they were used to identify errors or weaknesses in writing, to generate an initial draft of the Conclusion (Section 8), and to generate bibtex.
  • LLMs were used for retrieval and discovery. Specifically, they were used for finding related work.
  • LLMs were used for other purposes, such as the generation of the samples (described in Section 2) and streamlining code development.
Appendix B Additional figures
@@
Figure 3: Heatmaps of compliance performance. The left panels show the distribution of compliance ratings (in ‘confusion matrix’), comparing the median human expert with LLMs. The right panels show mean absolute error (MAE) across use cases and articles. Results are shown for Gemini 2.5 Pro (top) and as an average over all evaluated LLMs (bottom).
@@
Figure 4: Pareto frontier of model cost versus compliance agreement (Cohen’s $\kappa$). Each point represents a model, plotted by price (x-axis) and agreement with human expert ratings (y-axis). Pareto-efficient models are shown with red markers. Labels denote model names.
Table 3 presents the accuracy of LLMs in replicating human expert scores exactly and, where mismatches occur, the direction of error. Despite prompts designed to mitigate sycophancy and acquiescence bias, some models tended to assign higher compliance scores than human experts. Two of the worst models along this dimension were o3 mini and GPT-4o, with o3 mini strictly exceeding the median human expert score in 54.2% of excerpts, while only strictly falling below in 1.7% of excerpts. The model least prone to over-estimating the compliance level of excerpts was Gemini 2.5 Pro, which was also the model whose scores exactly matched those of the median human expert most frequently (for 60% of excerpts). Gemini 2.5 Flash achieved the best F1 score (0.913), meaning it excelled at correctly flagging compliant cases (score 4-5) and avoiding mislabelling non-compliant ones (score 1-3).
@@
Table 4 presents ablation results for GPT-4o, comparing the baseline model to three ablated versions: one without a tone prompt (see Appendix G), one with a harsh tone prompt, and one without access to the relevant AI Act text. The baseline tone prompt (“Your scores for both compliance and plausibility should be well-calibrated and objective. They should be rigorous but fair.”) achieves the best performance with respect to Cohen’s $\kappa_{w}$, Spearman’s $\rho$, and MAE.
@@
Table 5: Columns report agreement between LLM Saul-7B-Instruct (Colombo et al., 2024b) and Saul-54B-Instruct (Colombo et al., 2024a) and median human compliance scores across AIReg-Bench: quadratically weighted Cohen’s $\kappa_{w}$; Spearman’s $\rho$; Bias (mean signed difference, LLM-human); and MAE (mean absolute error). Outputs from Saul-7B-Instruct were resampled until they were provided in a parsable format.
Appendix C Design criteria
The design criteria for AIReg-Bench were as follows:

### v3

@@ 5.2 LLMs for legal, compliance, and AIR compliance tasks
Researchers have applied LLMs to a wide variety of legal tasks (Ma et al., 2024; Lai et al., 2024; Siino et al., 2025). This includes various compliance assessment tasks (Hassani, 2024; Bolton et al., 2025; Chen et al., 2024; Wang et al., 2025a), including AIR compliance assessments. For example, Makovec et al. (2024) input datasets, model cards, README files, or other AI project artifacts into a RAG-enhanced GPT-4 that accesses relevant portions of the AIA to predict the compliance level of the AI system depicted in the input. Davvetas et al. (2025) use a RAG-equipped LLM (mistral-small3.2) that takes, as input, certain features of an AI system (such as the type of AI system and the intended use) and outputs the risk-level of the AI system according to the AIA. Similarly, Kövári et al. (2025) use in-context learning and RAG to create a chatbot that can help users self-assess compliance with the AIA. Meanwhile, Li et al. (2025) test various LLMs’ ability to act as a rudimentary AIA compliance checker by accepting a hypothetical AI system as context and predicting whether it is prohibited by, permitted by, or out of scope of the AIA. A series of interlinked studies by Nokia Bell Labs employ LLMs to support AI practitioners in AIR compliance subtasks such as populating impact assessment reports (Bogucka et al., 2024a; Herdel et al., 2024; Bogucka et al., 2024b). Sovrano et al. (2025) use an LLM to help human-drafted technical documentation align with the requirements of AIA Article 11.
@@ 5.3 The EU AI Act
The AIA went into force in August 2024 (Lomas, 2024) and sets forth harmonized requirements for AI systems and models placed on the market or put into service in the EU (EU, 2024, Arts. 1-2). In laying out requirements for these AI systems, the AIA leverages a “risk-based” approach (Mahler, 2022), by which the exact requirements that apply to a system are a function of its perceived degree of risk. Here, the most demanding requirements are reserved for those AI systems deemed to be high-risk (EU, 2024, Art. 6). Such high-risk AI systems (HRAI) must satisfy a number of requirements (EU, 2024, Chap. III, Sec. 2). Among those, the requirements that we have made the focus of our benchmark dataset relate to risk management systems, data and data governance, record keeping, human oversight, as well as accuracy, robustness, and cybersecurity (EU, 2024, Art. 9, 10, 12, 14, 15).
5.4 Non-LLM algorithms that assess AIR compliance
It is important to distinguish the works in Section 5.2 as well as this work from those works that use algorithms other than LLMs, including but not limited to benchmark suites, to evaluate whether AI systems, including LLM-based systems, to comply with AI regulations (Prandi et al., 2025; Marino et al., 2024; Guldimann et al., 2024; Walke et al., 2023). While these works potentially present interesting accompaniments to the LLM-based approaches that this work aims to benchmark, this work does not seek to create a benchmark for these other methods.
@@ 6.1 Challenges of legal benchmarking
Compliance assessments are subjective. Complicating legal benchmarking is the fact that legal tasks often involve subjective judgments (Guha et al., 2024; Ma et al., 2023). Compliance, in particular, has been described as “not binary” (Wu and van Rooij, 2021). This was one motivation for our use of Likert scale annotations keyed to the “probability” of compliance rather than binary labels of “compliant” or “non-compliant.” The Likert scale is viewed as a reliable way to “transform subjective qualitative data into quantifiable metrics” (Koo and Yang, 2025) and has previously been used for benchmarking of LLMs in subjective realms (Bojić et al., 2025).
@@
Compliance assessments are a moving target. It has been said that compliance assessments require clear and specific guidelines, including relevant case law (Kilian et al., 2025; Schuett, 2024). But, in its present state, the AIA arguably lacks these. The text of the law has not been interpreted by courts (Yew et al., 2025). The obligations outlined in the AIA have yet to be clarified by accompanying technical standards (European Committee for Electrotechnical Standardization, 2024; Commission, 2022). They are also subject to ongoing amendments (EU, 2024, Art. 96). Accordingly, AIReg-Bench should be viewed as a snapshot of AIA conformity assessments in September 2025. It does not and cannot reflect developments occurring after this date.
@@ 6.2 Challenges of LLM-driven sample generation
Benchmark dataset samples should be representative of real-world data (Sourlos et al., 2024). Broadly speaking, there is evidence that LLMs (especially larger ones) can effectively produce synthetic samples satisfying this criteria (Maheshwari et al., 2024). More directly relevant here, it has been shown that LLMs can effectively generate (or improve) legal documents (Su et al., 2025; Lin and Cheng, 2024; Hemrajani, 2025; Gray et al., 2025), technical specifications (Xie et al., 2025), compliance documentation (Wang et al., 2025d; Hassani, 2024; Kumar and Roussinov, 2024), and even the type of technical documentation required under the AIA (Sovrano et al., 2025).
However, some are sceptical about the ability of LLMs to generate realistic outputs in these domains (Posner and Saran, 2025; Roberts, 2025; Shen et al., 2022). Critics point out that LLMs often lack domain-specific tacit knowledge, have difficulty maintaining coherent reasoning across extended contexts, and may hallucinate facts or references (Rasiah et al., 2024; Huang et al., 2024; Dahl et al., 2024; Magesh et al., 2025). Regarding legal tasks specifically, critics note how LLMs’ struggle to interpret legal terminology, grasp case context, and execute complex analyses, potentially resulting in errors (Wang et al., 2024; Roberts, 2025; Shen et al., 2022). With this in mind, a number of guardrails were put in place to enforce plausibility in our dataset samples and to create transparency around whether and where those guardrails fell short. For example:
  • The sample generation method was informed by the series of interviews with actual compliance assessment experts, described in Appendix C. We also consulted similar interviews performed by Li and Goel (2025) and the recommended protocols for manual compliance assessments for the AIA (Floridi et al., 2022; Thelisson and Verma, 2024; Lillo Campoy et al., 2024; Palumbo et al., 2025) and other AIR (The Institute of Internal Auditors, 2023; National Institute of Standards and Technology, 2023; Brogle et al., 2025).
@@ 6.3 Challenges of LLM-driven compliance analyses
Beyond the challenges of generating realistic documentation samples with LLMs, there may also be hurdles to using LLMs to perform legal analyses on text. Although some research efforts have found that LLMs match or exceed human accuracy when performing such analyses (Martin et al., 2024), others have questioned whether LLMs can perform this task effectively (Buckland, 2023; Doyle and Tucker, 2024; Li et al., 2024; Network, 2025; Mik, 2024). While the evidence presented in this paper is not definitive, we hope that our benchmark offers an initial step towards clarifying how well LLMs perform at compliance assessments, both in comparison to human experts and to one another. Our hope is that this will inspire other efforts to quantitatively evaluate the performance of LLMs at AIR compliance analyses and other legal tasks.
@@ 6.4 Risks of LLM-driven compliance analyses
As discussed in Section 1, there are potential benefits to LLM-driven AIR compliance analyses. However, it is important to point out that they could also carry risks. For example, given LLMs’ ongoing tendency to make errors, some argue that there may be dire consequences when lawyers place “blind faith in an LLM” (Moriarty, 2023) and that it may even represent a violation of professional ethics (Browning, 2024). Conversely, the developers of AI systems could try to “game” these LLMs, manipulating their technical documentation so as to achieve desirable compliance assessment outcomes. To help inform the conversation about these risks, and avoid generalization within that conversation, we believe it is invaluable to quantitatively evaluate the performance of LLMs for the task at hand, comparing their performance to human performance at that task in an evidence-based manner. Hence this benchmark.
@@ 7 Future Work
This benchmark serves as a starting point for tracking LLM progress at AIR compliance assessments, rather than as a finish line indicating readiness for deployment in legal practice. Moving towards that goal will require additional benchmarks that build upon AIReg-Bench. Suggested directions for extensions to AIReg-Bench are outlined below.
@@
Extension to further LLM-powered annotators. It would be valuable to extend our benchmarking in Section 4 to include more models: including fine-tuned legal LLMs, which some argue perform better at legal tasks (Fei et al., 2023; Dominguez‑Olmedo et al., 2024), as well as LLMs with tool-use (e.g., RAG or web search) (Makovec et al., 2024; Davvetas et al., 2025; Wang and Yuan, 2025). Though, while to our knowledge, no such models exist yet, there would also be value in fine-tuning LLMs for AIA or AIR compliance and then evaluating them using AIReg-Bench.
Extension to real-world documentation. AIReg-Bench consists of LLM-generated excerpts of technical documentation for the AIA. We relied on these LLM-generated excerpts, whose plausibility we manually verified, since public access to authentic technical documentation from real AI developers is limited, perhaps due to the confidentiality, legal privilege, or the relative nascency of the AIA. That said, we recognize the value, specifically as it relates to construct validity, of a benchmark built from real instances of AIA technical documentation. We therefore encourage any actors with access to such documentation to consider publishing (an anonymized version of) it and, furthermore, to contact us if we can assist with that process.
Extension to AIA technical standards. Like other EU product regulation, the AIA will utilize harmonized standards, i.e., “more concrete” (Siegmann and Anderljung, 2022) technical specifications prepared by the EU’s external standardization organizations (CEN, CENELEC and ETSI). Compliance with these specifications, which are still under development (European Committee for Electrotechnical Standardization, 2024; Commission, 2022), will “have the legal effect of establishing a presumption of conformity” with the AIA (Mazzini and Scalzo, 2023). Once issued, it would be important to pass these specifications, as context, into any AIA compliance assessment algorithm.
Extension to multi-turn interactions. Our interviews of AIR compliance professionals suggested that, in practice, compliance assessments often involve a long and complex dialogue between legal teams, technical staff, and regulators (see Appendix C). AIReg-Bench condenses the entire compliance assessment process into a single-turn interaction, based on a fixed set of synthetic artifacts. Such scoping and simplifications are common in benchmarking, though many scholars stress the need to shift towards more interactive modes of evaluating AI system capabilities (Ibrahim et al., 2024; Eriksson et al., 2025). Future benchmarks could build on AIReg-Bench by evaluating LLMs’ multi-turn ability to collaborate with human teams and contribute to complex legal dialogues (Kövári et al., 2025), rather than merely producing one-shot assessments.
@@ 8 Conclusion
Acknowledgements
We wish to acknowledge the Google Academic Research Awards for their support. Additionally, we wish to acknowledge the support of the OpenAI Researcher Access Program for providing subsidized API credits used in this research. We thank Bhagyashree Pancholy, Lara Groves, and Dr. Christoph Poetsch (TUV AI.Lab) for their insights during the early stages of this project. Additionally, we thank Kush Amerasinghe, Karl Muller, Sayam Kumar, Jonathan Marcotte, Ranjana Raghavan, Corey Smith, Maira Ata, and Ray Wang for assisting with annotation experiments.
Appendix A Additional figures
@@
Figure 3: Heatmaps of compliance performance. The left panels show the distribution of compliance ratings (in ‘confusion matrix’), comparing the median human expert with LLMs. The right panels show mean absolute error (MAE) across use cases and articles. Results are shown for Gemini 2.5 Pro (top) and as an average over all evaluated LLMs (bottom). Use1-8 reflect the intended uses from Table 8.
@@
Figure 4: Pareto frontier of model cost versus compliance agreement (Cohen’s $\kappa$). Each point represents a model, plotted by price (x-axis) and agreement with human expert ratings (y-axis). Pareto-efficient models are shown with red markers. Labels denote model names. We were not able to capture the cost of our particular implementation of Gemma, as API use is free but has a low restriction on output tokens.
Table 3 presents the accuracy of LLMs in replicating human expert scores exactly and, where mismatches occur, the direction of error. Despite prompts designed to mitigate sycophancy and acquiescence bias, some models tended to assign higher compliance scores than human experts. Two of the worst models in this regard were o3 mini and GPT-4o, with o3 mini strictly exceeding the median human expert score in 54.2% of excerpts, while only strictly falling below in 1.7%. The model least prone to over-estimating the compliance level of excerpts was Gemini 2.5 Pro, which was also the model whose scores exactly matched those of the median human expert most frequently (for 60% of excerpts). Gemini 2.5 Flash achieved the best F1 score (0.913), meaning it excelled at correctly flagging compliant cases (score 4-5) and avoiding mislabelling non-compliant ones (score 1-3).
@@
Table 4 presents ablation results for GPT-4o, comparing the baseline model to three ablated versions: one without a tone prompt (see Appendix F), one with a harsh tone prompt, and one without access to the relevant AI Act text. The baseline tone prompt (“Your scores for both compliance and plausibility should be well-calibrated and objective. They should be rigorous but fair.”) achieves the best performance with respect to Cohen’s $\kappa_{w}$, Spearman’s $\rho$, and MAE.
@@
Table 5: Experiments with Saul: Columns report agreement between LLM Saul-7B-Instruct (Colombo et al., 2024b) and Saul-54B-Instruct (Colombo et al., 2024a) and median human compliance scores across AIReg-Bench: quadratically weighted Cohen’s $\kappa_{w}$; Spearman’s $\rho$; Bias (mean signed difference, LLM-human); and MAE (mean absolute error). Outputs from Saul-7B-Instruct were resampled until they were provided in a parsable format.
GPT-5
GPT-4o
o3
o3 mini
Sonnet 4
Gemini 2.5 Pro
Gemini 2.5 Flash
Gemma 3
Grok 4
Grok 3 mini
GPT-5
GPT-4o 0.73
o3 0.70 0.57
o3 mini 0.55 0.87 0.45
Sonnet 4 0.84 0.75 0.74 0.55
Gemini 2.5 Pro 0.84 0.70 0.71 0.56 0.85
Gemini 2.5 Flash 0.73 0.67 0.75 0.51 0.77 0.74
Gemma 3 0.70 0.73 0.55 0.69 0.62 0.64 0.64
Grok 4 0.85 0.84 0.59 0.65 0.86 0.86 0.72 0.61
Grok 3 mini 0.68 0.92 0.55 0.77 0.78 0.67 0.70 0.78 0.82
Table 6: Pairwise Cohen’s Kappa scores. This table shows the results of performing pairwise Cohen’s Kappa scores of the models in our evaluation.
Table 7 summarizes the results of the alternative annotator test proposed by Calderon et al. (2025). The goal is to check whether replacing a human annotator with a given model would preserve the overall decisions. The winning rate captures the fraction of humans for whom the model would serve as an adequate substitute, while the average advantage probability reflects how often the model’s annotation is at least as good as a human’s across items. In our discrete label setup, a model wins on an item if its label is closer to the remaining annotators’ labels than the human’s.
Model Items Annotators Winning Rate Average Advantage Probability
GPT-5 60 3 0.6667 0.8500
GPT-4o 60 3 0.0000 0.7944
o3 60 3 0.0000 0.6556
o3 mini 60 3 0.0000 0.7556
Sonnet 4 60 3 0.0000 0.7333
Gemini 2.5 Pro 60 3 1.0000 0.9111
Gemini 2.5 Flash 60 3 0.0000 0.6722
Gemma 3 60 3 0.0000 0.6722
Grok 4 60 3 0.6667 0.8444
Grok 3 mini 60 3 0.0000 0.7500
Table 7: Alternative Annotator Test. This table shows the results of the test proposed by Calderon et al. (2025), which answers the question: if we replaced human annotators with another LLM, would the resulting labels stay the same? Here, winning rate is defined as the fraction of human annotators for whom the LLM passes the test (i.e., the model’s annotation is closer to the other human scores than the alternative annotation). Meanwhile, average advantage probability is defined as the average probability that the LLM’s annotation is at least as good as a human experts’s on each item. Our implementation reproduces the paper’s procedure as-is.
Use 1 Use 2 Use 3 Use 4 Use 5 Use 6 Use 7 Use 8 Total
Likert score 1 1 2 0 0 1 0 0 0 4
Likert score 2 6 5 4 1 6 1 3 4 30
Likert score 3 9 5 7 3 9 6 7 4 50
Likert score 4 11 17 11 16 16 19 15 17 122
Likert score 5 18 16 23 25 13 19 20 20 154
All scores 45 45 45 45 45 45 45 45 360
Table 8: AIReg-Bench dataset overview of human annotations for plausibility on a Likert scale. The uses correspond to, in order: traffic safety, gas delivery, education, exam proctoring, job hiring, job termination, emergency dispatch, and credit scoring.
Appendix B Design criteria
The design criteria for AIReg-Bench sample generation were as follows: